# Direct observation of the vanishing EELS cross section in graphene

### v1

@@ I.3 Intrinsic dielectric response of Graphene
In the top panel of Fig. 4, we show the ratio between the intensities of the $\pi$ and $\pi+\sigma$ plasmons as a function of momentum transfer $q$ for both the raw spectra and the dielectric response (obtained by dividing the spectra by $A$). Their spectral ratio $I_{\pi}/I_{\pi+\sigma}$ increases at lower momentum transfer due to the kinematic reshaping that lowers the intensity of the $\pi+\sigma$ plasmon more than that of the $\pi$ plasmon. Once the signal is divided by the reshape function $A$ and the intrinsic dielectric response is uncovered, the intensity ratio is instead approximately constant. According to Eq. (1), we deduce that in this regime the dielectric response is approximately converged to its optical value ($q=0$), while kinematic effects are still very sensitive to small momentum transfer variations.
Figure 4: Ratio between the intensities of the $\pi$ and $\pi+\sigma$ plasmons of graphene as a function of the momentum transfer $q$. The upper panel uses the experimental $q$. The lower panel uses $\tilde{q}$ inferred from the isotropic $\pi$ plasmon dispersion Wachsmuth et al. (2014); Guandalini et al. (2023a). Squares (circles) are $q$ and $\tilde{q}$ along the $\Gamma$M ($\Gamma$K) direction.
The finite accuracy of small momentum transfers is crucial for calculating $A$. We address the experimental accuracy in the lower panel of Fig. 4 by plotting the same intensity ratio with a momentum transfer $\tilde{q}$ calibrated on the well established isotropic linear $\pi$ plasmon dispersion Wachsmuth et al. (2014); Guandalini et al. (2023a). We note that a different momentum calibration does not just trivially change the $\tilde{q}$ values in the spectral intensity ratios, but it also alters the intensity ratios, as $A$ also changes. The qualitative picture is unchanged with the calibrated $\tilde{q}$, but the results are far less affected by experimental accuracy. The differences between $q$ and $\tilde{q}$ are below $\approx 0.02$-1, which is in line with the actual pixel size in the low magnification diffraction mode. $\tilde{q}$ has the potential to become a valuable tools for further instrumental developments regarding small angle TEM-EELS. $\tilde{q}$ can be determined from one single EEL spectrum of graphene and then the experimental $q$ can be calibrated against $\tilde{q}$.
@@ II Conclusions
III Supporting information
III.1 Experimental setup
Details on the sample preparation and specific conditions for EELS with an energy and momentum resolution (FWHM) of 45 meV and 0.02 Å-1, respectively, are given in Ref. Guandalini et al. (2023a). A sketch of our experimental setup is displayed in the left panel of the letter Fig. $1$. The JEOL3C2 TEM was operated at 30 kV, with a beam current of about 10 pA. The individual spectra are sums of 300 to 500 individual acquisitions with dwell times ranging between 1 ms and 1 s.
The resolution in $q$ depends on the electron wavenumber $k_{i}$, the half convergence angle ($\alpha$) of the incoming beam as well as the half collection angle ($\beta$) defined by the aperture in the diffraction plane and the camera length. With the 500 nm wide beam we achieve $q_{\alpha}=k_{i}\sin\alpha=0.006$-1 and $q_{\beta}=k_{i}\sin\beta=0.015$-1. The beam path stability amounts to an additional shared uncertainty of $q_{\vartheta}=k_{i}\sin\vartheta=0.004$-1. The resolution is thus given by
$\Delta q=k_{i}\left(\sqrt{\sin^{2}\alpha+\sin^{2}\beta}+\sin\vartheta\right)=0.02\ \text{\text{\AA}}^{-1}\ .$ (5)
In the left panel of Fig. 5 we plot the spatial and momentum resolution used in this work with the JEOL3C2 setup, alongside those obtained using other TEM setups Senga et al. (2019); Li et al. (2020); Hage et al. (2020); Senga et al. (2022); Hage et al. (2018); Hong et al. (2020); Guandalini et al. (2023a). The spot size determines the spatial resolution as well as the diffraction (Abbe) limit of the convergence angle $\alpha$, that could be attained only with perfect aberration correction. As shown in the left panel of Fig. 5, the achieved beam parameters $q_{\alpha}$ and $q_{\beta}$ are well above the Abbe limit of diffraction. For optimal performance the collection angle $\beta$ must be big enough for $\alpha$ and $\vartheta$. Smaller $\beta$ lose intensity without any meaningful effect in $\Delta q$. We did not include very asymmetric settings $\beta\ll\alpha$ with longer camera lengths (e.g. Ref. Midgley (1999)), as $q_{\beta}$ alone is not comparable to $\Delta q$. In the central panel of Fig. 5, we show a TEM image of the sample, along with a red mark with 500 nm diameter. This is the size of the collimated beam for the diffraction image in the right panel and the EELS measurements.
The zero-loss peak (ZLP) becomes more and more intense as the aperture approaches the optical limit. This challenges the possibility to observe the predicted quenching of the loss function. In order to lessen the impact of the ZLP, the profile of the direct beam in the diffraction plane must not hit the aperture. This should arguably always be easier to achieve with a pinhole-type aperture than with a slit.
The challenge with the very high momentum resolution and the according highly magnified diffraction planes is that the Bragg spots can not be recorded simultaneously, hence they cannot be used to establish the magnitude of $q$. The accuracy of the position in the diffraction plane is intrinsically limited to the pixel size (0.02 Å-1) in the overview diffraction pattern in the right panel of Fig. 5. The stability of the beam and reproducibility of the electron optics can be verified by checking for shifts when repeatedly zooming in and out.
Figure 5: Left: achieved resolutions in $q$ above the blue Abbe diffraction and orange Heisenberg uncertainty limits. Center: overview TEM micrograph. The size of the collimated beam is marked in red. Right: low magnification (x0.2) diffraction pattern.
III.2 EELS cross section of a 2D material at small momentum transfer
We consider a high energy electron incoming to a 2D material with momentum $k_{i}$ perpendicular to the layer and energy $E_{i}=\hbar^{2}k_{i}^{2}/2m$, where $m$ is the electron mass. Within scattering theory in the Born approximation Sturm (1993); Nazarov (2015), the (transmission) EELS differential cross section is given by
[…]
where $a_{0}$ is the Bohr radius, and $A^{u.c.}$ is the area of the 2D unit cell. The density-density response function of the system, defined as $\chi\equiv\delta\rho^{(1)}/\delta\phi^{ext}$, with $\phi^{ext}$ the perturbing scalar potential and $\rho^{(1)}$ the induced charge density, is Fourier transformed along the 2D periodic direction of the system in terms of the in-plane component of the momentum transfer $\mathbf{q}$. The out-of-plane momentum transfer $q_{z}$ cannot be interpreted in terms of crystalline quasi-momentum of the 2D system, that is defined only along the periodic directions, being instead formally accounted for by the double-integral along $z$ and the complex exponential in Eq. (6). In other words, $q_{z}$ is a good quantum number for the incident electron but not for the electrons of the system, while $q$ is a good quantum number for both kinds of electrons. If the system is composed by a sufficiently high amount of layers, the integrals in Eq. (6) give approximately the Fourier transforms along $z$ and $z^{\prime}$, and one recovers the bulk approximation for thick films.
Since we are interested in valence band excitations ($\hbar\omega\approx 1$-$20$ eV), with standard experimental setups ($E_{i}\approx 30$ keV, $k_{i}\approx 88.73$-1), $q_{z}$ ranges between $10^{-3}$ and $10^{-2}$-1 when approaching the optical limit $\mathbf{q}\rightarrow 0$. It follows that for atomically thin samples ($d\approx 1-5$) and in the low-momentum transfer regime, the complex exponential in Eq. (6) may be approximated to unity and the intrinsic dielectric response of the systems becomes independent of $q_{z}$. On the other hand, the presence of a finite $q_{z}$ in the denominator in front of the density-density response of Eq. (6) removes the diverging behaviour of the cross section when approaching the optical limit.
We are interested in the limit of small momentum transfers. If $qd\ll 1$, finite thickness effects are negligible and the system may be modeled with a dielectric sheet located at $z=0$ Sohier et al. (2015). In this limit, it is possible to define a 2D inverse dielectric function Sohier et al. (2015)
$\varepsilon^{-1}_{2D}(\mathbf{q},\omega)\equiv 1+v_{2D}(q)\tilde{\chi}(\mathbf{q},\omega)\ ,$ (7)
[…]
$\frac{d^{2}S}{d\Omega d\hbar\omega}(\mathbf{q},\omega)=\frac{8\varepsilon_{0}A^{u.c.}}{\pi^{2}e^{2}a_{0}^{2}}\frac{q}{(q^{2}+q_{z}^{2})^{2}}\mathrm{Im}\left[-\varepsilon^{-1}_{\mathrm{2D}}(\mathbf{q},\omega)\right]\ ,$ (8)
thus taking the form of a product between a purely kinematic factor $q/(q^{2}+q_{z}^{2})^{2}$ and the 2D loss function, in close analogy with the expression of HREELS cross section Mills (1975); Lambin et al. (1987); Hogan et al. (2009). In addition, the long-wavelength limit of the 2D dielectric function may be written as Cudazzo et al. (2011)
$\varepsilon_{\mathrm{2D}}(\mathbf{q}\to 0,\omega)\approx 1+i\frac{\tilde{\sigma}(\omega)e^{2}q}{2\varepsilon_{0}\omega}\ ,$ (9)
where $\tilde{\sigma}(\omega)\equiv\hat{\mathbf{q}}\cdot\bm{\sigma}_{\mathrm{2D}}(\omega)\cdot\hat{\mathbf{q}}$, and $\sigma^{ij}_{\mathrm{2D}}=\delta j^{(1),i}_{\mathrm{2D}}/\delta E^{j}$ is the 2D optical conductivity, relating the total electric field with the induced current density at $z=0$. Except for the case of acoustic plasmons, in the limit of small momentum transfers $\varepsilon_{\mathrm{2D}}(\mathbf{q}\to 0,\omega)$ linearly approaches unity (see Eq. 9), as shown in the static case for 2D semiconductors  (Qiu et al., 2016; Hüser et al., 2013; Guandalini et al., 2023b), and the imaginary part of $\varepsilon_{\mathrm{2D}}$ vanishes. From a physical perspective, the above limit may be understood as follows: in the case of two far apart test charges (i.e. in the long-wavelength limit), the 2D system is not able to screen the Coulomb interaction, as the vast majority of field lines bypass the material. The conductivity is instead finite, as finite current densities are allowed in the long-wavelength limit: the conductivity of graphene at zero momentum transfer is indeed the quantized conductance Tombros et al. (2011).
By inserting Eq. (9) into Eq. (8) the EELS cross section may be written as
$\frac{d^{2}S}{d\Omega d\hbar\omega}(|\mathbf{q}|\approx 0,\omega)=\frac{4A^{u.c.}}{\pi^{2}a_{0}^{2}}\frac{q^{2}}{(q^{2}+q_{z}^{2})^{2}}\frac{\mathrm{Re}\left[\tilde{\sigma}(\omega)\right]}{\omega}.$ (10)
Because of the overall kinematic prefactor $I_{kin}\equiv q^{2}/(q^{2}+q_{z}^{2})^{2}$, the cross section is quenched when the in-plane component $q$ of the momentum transfer $q_{\mathrm{3D}}$ approaches zero, even though the optical conductivity remains finite.
There is no possibility for the incoming electrons to excite a 2D system without also transferring a finite in-plane momentum $q$, as the intrinsic response of the system depends only on the latter. This feature is unique to low-dimensional systems, and is not captured by the bulk 3D approximation of the TEM-EELS cross section, that should apply only to thick enough films.
According to Eq. (1), the cross section quenching should be detected when $q\leq q_{z}$. As $q_{z}$ is of the order of 0.01 Å-1 in standard experimental setups, the momentum resolution should be at most of the same order of magnitude to reach the low-momentum transfer regime in experiments. On the other hand, $d\approx 3.5$  in freestanding graphene, implying that at $q\leq 0.1$ both conditions $q_{z}d\ll 1$ and $qd\ll 1$ are satisfied.
III.3 Optical properties and electron energy loss
The vanishing momentum transfer regime is also referred to as the optical limit, since photons transfer negligible momentum to the electrons. In the literature, it is generally accepted that in the long-wavelength limit ($q\to 0$) the EELS cross section provides the same information as optical absorption for 2D systems, as (finite-energy) collective excitations are quenched Stern (1967). The reflection and transmission coefficients of a 2D sheet for light polarized along one of its principal axes are
$\mathcal{R}(\omega)=\left|\frac{Z_{0}\tilde{\sigma}(\omega)}{2+Z_{0}\tilde{\sigma}(\omega)}\right|^{2}\ ,$ (11)
$\mathcal{T}(\omega)=\left|\frac{2}{2+Z_{0}\tilde{\sigma}(\omega)}\right|^{2}\ ,$ (12)
where $\tilde{\sigma}(\omega)$ is the optical ($q=0$) conductivity along the principal axis, and $Z_{0}=\sqrt{\nicefrac{{\mu_{0}}}{{\varepsilon_{0}}}}\approx 377\ \Omega$ is the vacuum impedance. The optical absorption coefficient $\mathcal{A}(\omega)$ may be obtained from the total energy conservation: $\mathcal{A}+\mathcal{R}+\mathcal{T}=1$ $\forall\omega$. Both optical absorption and EELS probe the same response function, the optical conductivity. Thus, the two experiments are linked by Eqs. (1) and (11)-(12). In addition, up to first order in $\tilde{\sigma}$, i.e., not in the resonant excitation regime, $\mathcal{R}(\omega)\approx 0$ and $\mathcal{A}(\omega)\approx 1-\mathcal{T}(\omega)\propto\mathrm{Re}[\tilde{\sigma}]$, as the EEL cross section.
In principle, therefore, optical conductivity information is embedded both in the absorption coefficient and in EEL spectra at low momentum transfer. However, the EELS cross section of Eq. (1) also depends on the kinematic prefactor $I_{kin}$, that is quenched in the very same regime of small momentum transfer. To access the intrinsic optical conductivity of the system, therefore, one has to appropriately account for kinematic effects in the low-$q$ EEL spectra.

### v3

@@ I.3 Intrinsic dielectric response of Graphene
In the (a) panel of Fig. 4, we show the ratio between the intensities of the $\pi$ and $\pi+\sigma$ plasmons as a function of momentum transfer $q$ for both the raw spectra and the dielectric response (obtained by dividing the spectra by $A$). Their spectral ratio $I_{\pi}/I_{\pi+\sigma}$ increases at lower momentum transfer due to the kinematic reshaping that lowers the intensity of the $\pi+\sigma$ plasmon more than that of the $\pi$ plasmon. Once the signal is divided by the reshape function $A$ and the intrinsic dielectric response is uncovered, the intensity ratio is instead approximately constant. According to Eq. (1), we deduce that in this regime the dielectric response is approximately converged to its optical value ($q=0$), while kinematic effects are still very sensitive to small momentum transfer variations.
Figure 4: Ratio between the intensities of the $\pi$ and $\pi+\sigma$ plasmons of graphene as a function of the momentum transfer $q$. The upper panel uses the experimental $q$. The lower panel uses $\tilde{q}$ inferred from the isotropic $\pi$ plasmon dispersion [44, 15]. Squares (circles) are $q$ and $\tilde{q}$ along the $\Gamma$M ($\Gamma$K) direction.
The finite accuracy of small momentum transfers is crucial for calculating $A$. We address the experimental accuracy in the (b) panel of Fig. 4 by plotting the same intensity ratio with a momentum transfer $\tilde{q}$ calibrated on the well established isotropic linear $\pi$ plasmon dispersion [44, 15]. We note that a different momentum calibration does not just trivially change the $\tilde{q}$ values in the spectral intensity ratios, but it also alters the intensity ratios, as $A$ also changes. The qualitative picture is unchanged with the calibrated $\tilde{q}$, but the results are far less affected by experimental accuracy. The differences between $q$ and $\tilde{q}$ are below $\approx 0.02$-1, which is in line with the actual pixel size in the low magnification diffraction mode. $\tilde{q}$ has the potential to become a valuable tools for further instrumental developments regarding small angle TEM-EELS. $\tilde{q}$ can be determined from one single EEL spectrum of graphene and then the experimental $q$ can be calibrated against $\tilde{q}$.